Answer the next six questions on the basis of the information in Table 1 which shows the short-run cost curves for a typical competitive firm. Table 1 AFC MO ATC AVCA $200.00 S50.00 S150.00 S50.00 2 112.50 37.5075.0025.00 85.00 35.00 50.00 30.00 4 73.75 36.2537.5040.00 30.00 55.00 6 70.00 45.0025.0070.00 7286 51.4321.43 90.00 78.13 59.38 18.75 115.00 30 030 3304 30 0 5 70.00 40.00 3 60 5 605 60 6 60 0 75 6 75 6 75 775 6 95 795 795 95 7 9 85.56 68.89 0 95.0080.00I 16.67 145.00 15.00 180.00 Now assume there are 600 identical firms in this competitive industry. That is, there are 600 firms, each of which has the same cost data as the single firm in Table 1. Table 2 lists the market demand curve for this product. . Referring to the data in Table 1, if the product price was $70, in the short run, the firm would: A. leave the industry B. produce 10 units and incur a loss of $250. C. produce 5 units and only eam normal profits. D. produce 6 units and earn economic profits equal to SO E. shut down and incur a loss of $150. antity Demanded 5,975 2. According to Table 1, if the product price was S120, 5,125 in the short run, the profit-maximizing firm would: A. shut down and incur a loss of $150 B. produce 6 units and earn economic profits of $300. C. produce 10 units and earn economic profits of $250. D. produce 8 units and carn normal profits of S335 E. produce 8 units and 150 earn economic profits of S335. 6. Given the data in Tables 1 and 2, the short-run equilibrium price will be: 3. According to Table 1, if the market price was $60, in the short run, the profit-maximizing competitive A. $20 B. S30 firm would: A. produce 6 units and incur a loss of $60. B. produce 10 units and incur a loss of $350. D. $60. E S75. F. $95 C. produce 4 units and incur a loss of $55 D. produce 5 units and incur a loss of S50. E. shut down and incur a loss of $150. H $150 4. Referring to Table 1, if the product price was S30 a unit, in the short run, the profit-maximizing firm Questions 7-II are on the other side. would A. produce 3 units and incur a loss of $165. B. shut down and incur a loss of $150. C. produce 10 units and incur a loss of $650 D. produce 6 units and incur a loss of $240. E. produce 5 units and incur a loss of S200. 5. Based on the information in Table 1, which of the following represents the profit-maximizing, competitive firms short-run supply curve?
Answer #1

1. The correct answer is: d)

Reason: At profit maximizing level, P = MC

Here, this condition is satisfied at 6 unit of output where P = MC = 70.

Economic profit = (P - ATC) *Q = (70-70)*6 =0.

P.S.: as HOMEWORKLIB's policy if nothing is mentioned then only the first quwquest is to be answered

Thanks!